OpenStudy (cherie_magee):

Solve the following system by the comparison method. 2p + q = 1 9p + 3q + 3 = 0 What is the solution set? {(-2, 5)} {(2, -5)} {(-2, -5)}

OpenStudy (anonymous):

what is the "comparison" method?

OpenStudy (cherie_magee):

i really don't know :( @satellite73

OpenStudy (anonymous):

plug in the numbers and see if they work, that is probably easiest

OpenStudy (anonymous):

or else just solve \[2p + q = 1 \\9p + 3q + 3 = 0\] the first one is the same as \[q=1-2p\] substitute in to the second and get \[9p+3(1-2p)+3=0\] and solve for \(p\)

OpenStudy (anonymous):

you got this?

OpenStudy (cherie_magee):

yeah i got it thank ya @satellite73

OpenStudy (anonymous):

yw